A wheel has 38 numbers, 1, 2, 3, . . . , 38. A player picks a number and bets $20. The wheel is spun, and if the player􏰀s number results, he or she is paid $700 (and gets to keep the $20 bet). If another number results, the house keeps the $20. Suppose a typical player bets on 60 spins per hour. What is the expected net gain per hour for the house on each player?
Answer by Boreal(15235)

E(x)=(1/38)$720-20(37/38)
=18.95-19.47=-0.52 (don't round until end, however)
Multiply by 60 spins, and the expected lost is $31.58 per hour (round at end)